OpenStudy (anonymous):

Find the area under the curves. y=4sin(x) y=2cos(x) x=0 and .4pi Please help!

OpenStudy (anonymous):

Set-up an integral like \[\int\limits_{0}^{4\pi}4\sin x dx\] for each one, or use the fnInt feature on your calculator with fnInt(4*sin(x),x,0,4*pi) to spit out your answer. The second one is very similar.

OpenStudy (anonymous):

and then add them?

OpenStudy (anonymous):

Is it between the curves or the area under each curve?

OpenStudy (anonymous):

Find the area of the region enclosed between

OpenStudy (anonymous):

Ah, gotcha. Okay, so you would \[\int\limits_{0}^{.4\pi}[2\cos(x) - 4\sin (x)]dx\]. This way you subtract the larger function at x=0 by the smaller one there.

OpenStudy (anonymous):

I tried that earlier and it didn't work! :/

OpenStudy (anonymous):

Check your limits of integration, then. Is it really .4pi or is it 4pi? The graph crisscrosses a lot, so the integral above would only work if .4pi is the first intersection point, and according to my TI-86, it is not. Thus, I recommend splitting up the integral into two pieces: one will go from 0 to the intersection point of x=.464 as set up above, the other will go from .464 to .4pi with the order of the two functions switched.

OpenStudy (anonymous):

It definitely is .4pi. When you mean switch you mean 4sin(x)-2cos(x)?

OpenStudy (anonymous):

Yes, exactly, because from x=.464 to .4pi, 4sin(x) attains greater values than 2cos(x).